OpenStudy (anonymous):

A professional golfer teed off at the 13th hole of the U.S. Open. It took 0.0020 seconds for the head of her club to travel the last 10 centimeters (cm) before the club hit the ball. The following formula relates the velocity to the distance and time of the club¿s motion, where V is the velocity of the club in kilometers per hour (km/h), d is the distance the club travels in centimeters (cm), and t is the time in seconds (s): What is the velocity, in km/h, of her club during this swing? A. 18 B. 72 C. 180 D. 720

OpenStudy (anonymous):

Plug the numbers into the formula which they appeared to have given you.

OpenStudy (anonymous):

d=v*t*\[\frac{ 1000 }{ 36 }\]

OpenStudy (anonymous):

Okay, isolate the V. Can you do that?

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

Okay. Well, if we wanted to get V by itself, you would want to get rid of the t, right?

OpenStudy (anonymous):

So what's the inverse (opposite) operation of multiplication?

OpenStudy (anonymous):

ummmm i seriously have no clue how to do it

OpenStudy (anonymous):

I'm guiding you. Just answer the question.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

The inverse operation of + is -. The inverse operation of multiplication, is division. Remember this, it'll help you isolate stuff.

OpenStudy (anonymous):

SO NOW. When you divide both sides by t, you'll isolate V, y'see?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Okay, now you have the equation V = d/t, right?

OpenStudy (anonymous):

So plug in the distance and the time.

OpenStudy (anonymous):

i gtg i surley gonna guess maybe next time ill have time for the study part

OpenStudy (anonymous):

....well I'll finish my explanation I suppose. Even after you plug in the numbers, it really won't make sense, because they're asking for the answer in km/h, and the plug-in is in cm/sec.

OpenStudy (anonymous):

thnxs i guess

OpenStudy (anonymous):

Sorry I can't actually teach you everything that you need to know. I hope you go to extra help or whatever else they offer you at your school so that you can get these basics down, 'cuz if you did, you'd be able to do them so quickly, without having to turn to openstudy for all your questions. People aren't going to just hand you answers. You're going to have to learn how to do these things. I wish you luck and I hope you do well....

OpenStudy (mathstudent55):

\[v = \frac{ d }{ t } = \frac{ 10 \space cm }{ 0.0020 \space s } \times \frac{ 1 \space m }{ 100 \space cm } \times \frac{ 1 \space km }{ 1000 \space m } \times \frac{ 60 \space s }{ 1 \space mn } \times \frac{ 60 \space mn }{ 1 \space h } = 180 \space km/h\]

OpenStudy (mathstudent55):

The reason we can multiply the first fraction by all the others is that all other fractions are equal to 1. Since 1 m = 100 cm, then (1 m)/(100 cm) = 1. Since 1 km = 1000 m, then (1 km)/(1000 m = 1). Since 60 s = 1 mn, then (60 s)/(1 mn) = 1 Since 60 mn = 1 h, then (60 mn)/(1 h) = 1 All the above are conversion factors. They are all equal to the number 1 written in forms that allow us to convert units from cm/s to km/h.